# Initial Guessing Bias: How Untrained Networks Favor Some Classes

### v3

@@ I.1 Experiments on real data
In this section, we show some empirical evidence of IGB on combinations of models and data which are not covered by our theory, such as CNN architectures and the MNIST, CIFAR10 and CIFAR100 datasets. [^18] MNIST is a particularly insightful dataset for IGB, because the correlations between pixels are high. This is opposite to what we cover with our theory, which is based on i.i.d. inputs. In App. I.2, we will also include experiments on additional architectures (ResNet, MLP-mixer and Vision Transformers). The experiments we will discuss do not presume to show in a complete and exhaustive way the range of models and cases in which IGB can occur; this question is outside the scope of our work. Rather, they are some didactic examples intended to show how, beyond the assumptions used in our analysis to quantitatively treat the phenomenon, IGB also emerges in a broader context. We first present and discuss the experiments. For technical details for the reproducibility we refer the reader to App. I.3.
@@ I.1.1 A prototype for highly correlated data: MNIST
Figure I.1: Comparison of $p_{f_{0}}\left({x}\right)$ between two untrained neural networks fed with MNIST dataset; the ten classes of the dataset were merged in two macro groups: even and odd. The two architectures employed in the comparison differ in the choice of activation function and pooling. For these simulations we used the E&O dataset on the model CNN-B. (see App. I.3 for more details). On the right, the same plot in a logarithmic scale to better visualize the differences in the low-density regions.
@@
MNIST constitutes a good candidate to represent a scenario of strong correlations in the data. In fact, in this dataset the images are characterized by a small percentage of nonzero pixels; this leads to the formation of areas, within the same image, characterized by similar values and therefore strongly correlated. In addition, image areas are strongly correlated between different elements. For example, we are unlikely to observe the writing of a number on the edges of an image; therefore, different images will have similar values of pixels along the edges. When considering images of the same class, this correlation is even more pronounced. From the MNIST dataset, we defined two macro-classes, even and odd, by merging the 10 starting classes according to their parity. We then propagated the binary dataset, obtained in this way, through two untrained convolutional neural networks. The difference between the two networks lies in the choice of activation and pooling function employed. In particular the dataset is propagated through CNN-B (network details in App. I.3). The CNN-B model is a CNN where the last convolutional layer is directly connected to the outputs. We choose it this way, because, as earlier shown, IGB cannot arise without hidden layers. This indicates that the observation of IGB is caused by the convolutional layers or, to put it more precisely, it emerges from the propagation of structured data through the convolutional layers.
@@ I.1.2 A further example: CIFAR10
Figure I.2: Comparison of $p_{f_{0}}\left({x}\right)$ between two untrained neural networks fed with CIFAR10 dataset (with 10 classes). The two architectures employed in the comparison differ in the choice of activation function and pooling. In the absence of IGB, the distributions would concentrate around the vertical line. For these simulations we used the dataset C10 on the model CNN-B described below (see App. I.3). On the right, the same plot in a logarithmic scale to better visualize the differences in the low-density regions.
@@ I.1.3 High cardinality dataset: Cifar100
Figure I.3: CNN-A (ReLU+MaxPool case) with C100 in this case ($N_{C}=100$). The peak at $f_{0}=\frac{1}{N_{C}}$, corresponding to the No IGB case (in the absence of IGB, the distribution should concentrate around this value), is reported as a reference. The plot on the right contains the same data as that on the left, but has a logarithmic $x$ axis. we added a regularizer $\epsilon=1e-9$ to the value of $f_{0}$ in order to show the peak at $f_{0}=0$.
Figure I.4: Right: Comparison of the trend of $\max_{i}\{f_{i}\}$ with that of accuracy during the learning dynamics, varying with the level of guessing bias at initialization (IGB). The curves show a consistent pattern with the diagram on the shown in Fig. 2. The simulations were conducted on a ResNet (top) and Vision Transformer (bottom) using a binary dataset (dogs vs cats from CIFAR) as input; more details on the setting are provided in App. I.3.
Figure I.5: Right: Comparison of the trend of $\max_{i}\{f_{i}\}$ with that of accuracy during the learning dynamics, varying with the level of guessing bias at initialization (IGB). The curves show a consistent pattern with the diagram on the shown in Fig. 2. The simulations were conducted on different architectures (MLP- mixer (top), ResNet (middle) and Vision Transformer (bottom)) using a multi-class dataset (C10) as input; more details on the setting are provided in App. I.3.
@@ I.2 Experiments on other architectures & effects on the training dynamics
I.3 Reproducibility
Here we provide technical details about the experiments, to allow for reproducibility. The code used for the experiments presented in this work are available at https://anonymous.4open.science/r/Algorithms-82EC/README.md.
@@ I.3 Reproducibility
  • CIFAR100 (C100): We use CIFAR100 (https://www.cs.toronto.edu/~kriz/cifar.html) [Krizhevsky et al., 2009] as an example of high cardinality dataset, i.e. a dataset with a big number of classes
@@
  • MLP-mixer: We propose the MLP-mixer (MLP-mix) introduced in [Tolstikhin et al., 2021] as an example of a more advanced MLP architecture. As a basis for the proposed simulations, we will use the architectural design documented in https://github.com/omihub777/MLP-Mixer-CIFAR/blob/main/README.md.
  • ResNet: We propose ResNet34 (ResNet), introduced in He et al. [2016] as an example of an architecture equipped with skip connections. As a basis for the proposed simulations, we will use the architectural design documented in https://www.kaggle.com/code/kmldas/cifar10-resnet-90-accuracy-less-than-5-min.
  • Vision Transformer: We propose ViT (ViT), introduced in Dosovitskiy et al. [2020] as an example of an architecture equipped with multi-head attention mechanism. As a basis for the proposed simulations, we will use the architectural design documented in https://github.com/tintn/vision-transformer-from-scratch/blob/main/vit.py.

### v4

@@ I.1 Experiments on real data
In this section, we show some empirical evidence of IGB on combinations of models and data which are not covered by our theory, such as CNN architectures and the MNIST, CIFAR10 and CIFAR100 datasets. [^20] MNIST is a particularly insightful dataset for IGB, because the correlations between pixels are high. This is opposite to what we cover with our theory, which is based on i.i.d. inputs. In App. I.2, we will also include experiments on additional architectures (ResNet, MLP-mixer and Vision Transformers). The experiments we will discuss do not presume to show in a complete and exhaustive way the range of models and cases in which IGB can occur; this question is outside the scope of our work. Rather, they are some didactic examples intended to show how, beyond the assumptions used in our analysis to quantitatively treat the phenomenon, IGB also emerges in a broader context. We first present and discuss the experiments. For technical details for the reproducibility we refer the reader to App. I.4.
@@ I.1.1 A prototype for highly correlated data: MNIST
Figure I.1: Comparison of the distribution $p_{f_{0}}\left({x}\right)$ between two untrained neural networks fed with MNIST dataset; the ten classes of the dataset were merged in two macro groups: even and odd. The two architectures employed in the comparison differ in the choice of activation function and pooling. For these simulations we used the E&O dataset on the model CNN-B. (see App. I.4 for more details). On the right, the same plot in a logarithmic scale to better visualize the differences in the low-density regions.
@@
MNIST constitutes a good candidate to represent a scenario of strong correlations in the data. In fact, in this dataset the images are characterized by a small percentage of nonzero pixels; this leads to the formation of areas, within the same image, characterized by similar values and therefore strongly correlated. In addition, image areas are strongly correlated between different elements. For example, we are unlikely to observe the writing of a number on the edges of an image; therefore, different images will have similar values of pixels along the edges. When considering images of the same class, this correlation is even more pronounced. From the MNIST dataset, we defined two macro-classes, even and odd, by merging the 10 starting classes according to their parity. We then propagated the binary dataset, obtained in this way, through two untrained convolutional neural networks. The difference between the two networks lies in the choice of activation and pooling function employed. In particular the dataset is propagated through CNN-B (network details in App. I.4). The CNN-B model is a CNN where the last convolutional layer is directly connected to the outputs. We choose it this way, because, as earlier shown, IGB cannot arise without hidden layers. This indicates that the observation of IGB is caused by the convolutional layers or, to put it more precisely, it emerges from the propagation of structured data through the convolutional layers.
@@ I.1.2 A further example: CIFAR10
Figure I.2: Comparison of the distribution $p_{f_{0}}\left({x}\right)$ between two untrained neural networks fed with CIFAR10 dataset (with 10 classes). The two architectures employed in the comparison differ in the choice of activation function and pooling. In the absence of IGB, the distributions would concentrate around the vertical line. For these simulations we used the dataset C10 on the model CNN-B described below (see App. I.4). On the right, the same plot in a logarithmic scale to better visualize the differences in the low-density regions.
@@ I.1.3 High cardinality dataset: Cifar100
Figure I.3: CNN-A (ReLU + max pool case) with C100 in this case ($N_{C}=100$). The peak at $f_{0}=\frac{1}{N_{C}}$, corresponding to the No IGB case (in the absence of IGB, the distribution should concentrate around this value), is reported as a reference. The plot on the right contains the same data as that on the left, but has a logarithmic $x$ axis. We added a small numer, $\epsilon=1e-9$, to the value of $f_{0}$ in order to show the peak at $f_{0}=0$.
Figure I.4: Right: Comparison of the trend of $\max_{i}\{f_{i}\}$ with that of accuracy during the learning dynamics, varying with the level of guessing bias at initialization (IGB). The curves show a consistent pattern with the diagram on the shown in Fig. 2. The simulations were conducted on a ResNet (top) and Vision Transformer (bottom) using a binary dataset (dogs vs cats from CIFAR) as input; more details on the setting are provided in App. I.4.
Figure I.5: Right: Comparison of the trend of $\max_{i}\{f_{i}\}$ with that of accuracy during the learning dynamics, varying with the level of guessing bias at initialization (IGB). The curves show a consistent pattern with the diagram on the shown in Fig. 2. The simulations were conducted on different architectures (MLP- mixer (top), ResNet (middle) and Vision Transformer (bottom)) using a multi-class dataset (C10) as input; more details on the setting are provided in App. I.4.
@@ I.2 Experiments on other architectures & effects on the training dynamics
I.3 IGB in Pre-Trained Models
Thus far, the analysis and experiments presented have focused on randomly initialized DNNs. While the results in App. I.2 provide valuable insights into the effects of IGB on the learning dynamics of untrained networks, we now demonstrate that even pre-trained models exhibit the presence of IGB.
Transfer learning is becoming an increasingly prevalent paradigm [Han et al., 2021], where large models are pre-trained on extensive datasets and then fine-tuned for specific, different tasks. In practice, before initiating the fine-tuning phase, the final layer, referred to as the head, is replaced by untrained layer(s). This replacement ensures compatibility with the new task, such as modifying the number of output nodes in the classifier (i.e., the last fully connected layer) to match the number of classes in the new task.
While the head could, in principle, consist solely of the classifier, other configurations with more complex structures for the head are possible [Ren et al., 2023]. Using a more articulated head is a common practice, especially in situations where the head is the only part of the network trained during fine-tuning. Approaches that avoid full-model fine-tuning have recently received more attention, as they offer reduced computational costs and address privacy concerns [Xiao et al., 2023], while also ensuring better performance in the presence of out-of-distribution data with large distribution shifts [Kumar et al., 2022].
Fig. I.6 shows the presence of IGB on pre-trained models at the beginning of the fine-tuning phase. The experiments are conducted on different architectures; in each of them the last fully connected layer (classifier) is replaced to match the new number of classes and reinitialized.
Figure I.6: The distribution $p_{f_{0}}\left({x}\right)$ on pre-trained models. From left to right are shown PT-ResNet, PT-SwinT-FC, PT-ENetV2-FC on a binary dataset (Cats and Dogs from C10).
Interestingly, using a deeper untrained structure as head leads to an amplification of IGB as shown in Fig.I.7; the same architectures from experiments in Fig. I.6 are considered with the only difference that the fully connected head is now replaced by an untrained MLP.
Figure I.7: The distribution $p_{f_{0}}\left({x}\right)$ on pre-trained models. From left to right are shown PT-SwinT-MLP, PT-ENetV2-MLP on a binary dataset (Cats and Dogs from C10)
I.4 Reproducibility
Here we provide technical details about the experiments, to allow for reproducibility. The code used for the experiments presented in this work are available at https://github.com/EmanueleFrancazi/IGB-Algorithms.
@@ I.4 Reproducibility
  • CIFAR100 (C100): We use CIFAR100 (https://www.cs.toronto.edu/~kriz/cifar.html) [Krizhevsky et al., 2009] as an example of high cardinality dataset, i.e. a dataset with a big number of classes.
@@
  • MLP-mixer: We propose the MLP-mixer (MLP-mix) introduced in [Tolstikhin et al., 2021] as an example of a more advanced MLP architecture. We use the architectural design documented in https://github.com/omihub777/MLP-Mixer-CIFAR/blob/main/README.md.
  • ResNet: We propose ResNet34 (ResNet), introduced in He et al. [2016] as an example of an architecture equipped with skip connections. We use the architectural design documented in https://www.kaggle.com/code/kmldas/cifar10-resnet-90-accuracy-less-than-5-min.
  • Vision Transformer: We propose ViT (ViT), introduced in Dosovitskiy et al. [2020] as an example of an architecture equipped with multi-head attention mechanism. We use the architectural design documented in https://github.com/tintn/vision-transformer-from-scratch/blob/main/vit.py.
  • Pre-Trained Resnet: ResNet50 (PT-ResNet) pretrained on Imagenet [Deng et al., 2009] (for more details see https://pytorch.org/vision/stable/models/generated/torchvision.models.resnet50.html#torchvision.models.resnet50). The head is substituted by an untrained classifier (fully connected layer) randomly initialized with Kaiming normal weights and null bias.
  • Pre-Trained Swin Transformer: swin tiny architecture pretrained on Imagenet [Deng et al., 2009] (for more details see https://pytorch.org/vision/main/models/generated/torchvision.models.swin_t.html#torchvision.models.swin_t
    • PT-SwinT-FC: The classifier (last fully connected layer) is randomly re-initialized with Kaiming normal weights and null bias.
    • PT-SwinT-MLP: The classifier is substituted by a multi-layer perceptron (5 layers) with ReLU activation function randomly initialized with Kaiming normal weights and null bias.
  • Pre-Trained EfficientNet: EfficientNetV2-S architecture pretrained on Imagenet [Deng et al., 2009] (for more details see https://pytorch.org/vision/main/models/generated/torchvision.models.efficientnet_v2_s.html#torchvision.models.efficientnet_v2_s )
    • PT-ENetV2-FC: The classifier (last fully connected layer) is randomly re-initialized with Kaiming normal weights and null bias.
    • PT-ENetV2-MLP: The classifier is substituted by a multi-layer perceptron (5 layers) with ReLU activation function randomly initialized with Kaiming normal weights and null bias.
@@
By informing model selection, data preparation and initial conditions, our results can improve the training of machine learning models. Better-performing machine learning models allow to better address wide ranges of problems, but can also be adapted for potentially harmful applications [Hutson, 2021, Qadeer and Millar, 2021].